# T-JEPA: Augmentation-Free Self-Supervised Learning for Tabular Data

### v1

@@ 4.2 Results
Binary and Multi-Class Classification
As displayed in Table 1, for the classification tasks, T-JEPA pre-training improved the performance of all evaluated models except on the Higgs (HI) dataset where only the performance of ResNet improves when augmented by T-JEPA. On average T-JEPA increased the performance of MLP by 6.86.8% excluding HI and $0.7\%$ otherwise, DCNv2 by $7.5\%$ and $1.4\%$, ResNet by $7.1\%$ and $6.4\%$, AutoInt by $5.8\%$ and $0.32\%$ and FT-Transformer by $3.6\%$ and $1.1\%$. Interestingly, despite their design advantages for tabular data, recent attention-based models like FT-Transformer and AutoInt, though improved by T-JEPA, did not surpass architectures like ResNet and MLP, which exhibited the most significant gains with T-JEPA. This outcome suggests that T-JEPA may complement the inductive biases of more traditional architectures like ResNet and MLP, which, despite their simpler design, are better equipped to leverage the feature representations learned through T-JEPA. These findings underscore the importance of considering both model architecture and pre-training strategies when optimizing for tabular data tasks, highlighting T-JEPA’s potential as a tool for enhancing model performance. Finally, we observe that for classification tasks ResNet+T-JEPA appears as a strong candidate as it outperforms all other methods including GBDTs, which is in line with Gorishniy et al. (2021b) and Zabërgja et al. (2024).
Regression
In regression tasks, T-JEPA induces a consistent decrease of the RMSE on the California Housing (CA) dataset for all tested models, indicating improved performance. For the MLP model, the MSE improved by $9.7\%$, DCNv2 by $16.3\%$, ResNet by $17.9\%$, AutoInt by $11.7\%$ and FT-Transformer by $8.6\%$. This indicates that T-JEPA pre-training benefits the learning process for regression tasks.
Compute
T-JEPA pre-training consistently improves performance across a variety of deep learning models, demonstrating its model-agnostic benefits while also requiring moderate compute and training time as displayed in Table 5 in appendix B.
@@ 5.1 Representation space evaluation
5.1.1 Metrics
To assess the characteristics of the generated representation space, we rely on three metrics: Kullback-Leibler divergence, uniformity score and feature variances. These metrics provide insights into the distribution, compactness and separability of the embedded features.
Kullback-Leibler divergence
$D_{\text{KL}}(P\parallel Q)=\sum_{x}P(x)\log\frac{P(x)}{Q(x)}.$ (6)
Uniformity
We rely on the uniformity score (Wang and Isola, 2020) to evaluate the preservation of maximal information within the feature distribution. This score leverages the Gaussian potential kernel $G_{t}:\mathcal{S}^{d}\times\mathcal{S}^{d}\rightarrow\mathbb{R}_{+}$, defined as $G_{t}(u,v)\triangleq e^{-t\|u-v\|^{2}_{2}},\;t>0$. The uniformity score is defined as the logarithm of the average pairwise Gaussian potential, formally:
Feature variance
Finally, we also include intra-feature variance as a measure of the variability within individual features, and inter-feature variance to quantify the variability of the hidden dimension across different features. More precisely:
$\begin{array}[]{cc}\sigma_{\text{intra}}^{2}=\frac{1}{d}\sum_{i=1}^{d}\text{Var}(\mathbf{h}_{i,:}),&\sigma_{\text{inter}}^{2}=\frac{1}{h}\sum_{j=1}^{h}\text{Var}(\mathbf{h}_{:,j}),\end{array}$ (8)
where $\mathbf{h}\in\mathbb{R}^{d\times h}$ represents the embedding space with $d$ features and $h$ hidden dimensions.
5.1.2 Embedding space characterization
@@
$\sigma_{\text{intra}}^{2}$ $1.00$ $1.03$ $1.08$ $1.12$ $1.14$ $1.17$ $1.17$
$\sigma_{\text{inter}}^{2}$ $0.65$ $0.16$ $0.14$ $0.09$ $0.08$ $0.07$ $0.07$
@@
To assess whether T-JEPA generates representations uniformly spanned across the embedding space, we display in Figure 2 the obtained sample representations for the JA dataset at different training epochs. We randomly sample 50,000 points and rely on the PaCMAP dimensionality reduction technique to reduce their representations in the embedding space to two dimensions for visualization purposes. We observe in Figure 2 that T-JEPA effectively utilizes the representation space, evolving from an initial collapsed distribution towards a more structured arrangement as the training progresses. This behavior is indicative of the model’s capacity to learn distinct and meaningful representations, which are crucial for downstream tasks. We also display in Table 2 how the intra-feature and inter-feature variances ($\sigma_{\text{intra}}^{2}$, $\sigma_{\text{inter}}^{2}$), uniformity score, KL divergence ($D_{\text{KL}}$), and the euclidean distance ($\|\cdot\|_{2}$) vary during training of. At each epoch, we randomly select 50,000 sample representations from the JA dataset and compute these metrics. For both KL-Divergence and euclidean distance, we report the average pairwise KL-Divergence and euclidean distance for the 50,000 samples previously selected. We observe a stable intra-feature variance that indicates consistent information extraction, while the decreasing inter-feature variance shows improved feature specialization. The rising KL divergence and euclidean distance demonstrate that the model effectively avoids representation collapse, leading to more meaningful and discriminative embeddings. Finally, increasing uniformity across epochs is in line with the representations displayed in Figure 2 and demonstrates better separability.
@@ 5.2 Representation Collapse
Figure 3: Training regime of Joint-Embedding Predictive Architectures on tabular (JA) and image (ImageNet-1K) data. We display on the right a randomly selected sample’s representations for each critical part of the training process, where (a) describes the initial random initialization, (b) the collapsed equilibrium, (c) the regularization effect pushing the weights outside of the collapsed equilibrium, and (d) the convergence.
@@
We observe that JEPA-based models undergo an initial representation collapse due to the EMA relation between the weights of the context and label encoders. Indeed, we retrained from scratch I-JEPA (Assran et al., 2023) on ImageNet-1K using their official implementation and hyperparameters [^1] For computational purposes, we reduced the batch size to $16$ and kept unchanged the rest of the hyperparameters. and observed a similar training regime as for T-JEPA.
@@
First, the loss collapses close to 0 after a few iterations; then, regularization starts pushing the model’s weights toward a non-trivial equilibrium. We display in Figure 3 the training regimes of both T-JEPA and I-JEPA.
@@
While other JEPA methods do not appear to require further regularization, we observe that appending a regularization token [REG] to the sample’s representations is instrumental in escaping training collapse regimes. As displayed in Figure 4, when training T-JEPA on the Jannis dataset without including any regularization token, the optimization process does not manage to escape the initial collapse. On the contrary, when including one or more tokens, we can escape this initial collapse and push the weights towards a non-trivial equilibrium.
@@ 5.3 Comparison with Gradient Boosted Decision Trees
To try and understand why T-JEPA enabled some approaches to match the performance of GBDT, we investigated whether high-variance features in the latent space correlate with feature importance for downstream tasks. The embedding variance was computed as the standard deviation of each feature’s values across hidden dimensions, normalized by the dimension-wise mean. We ranked features by their average embedding variance across all samples and compared these rankings to those generated by traditional supervised methods, including XGBoost and permutation importance, using Kendall’s $\tau$ for rank similarity. Overall, our analysis reveals a strong correlation between high-variance features in the T-JEPA embeddings and those identified as important by supervised methods, despite T-JEPA being trained without target labels. Figure 5 provides a detailed comparison. The embedding variance ranking ($\sigma_{\text{embed}}$) shows significant correlation with XGBoost ($\tau=0.44$, with $p\text{-value}=1.73e^{-6}$). A random generated rank is provided for comparison. These findings suggest that T-JEPA’s self-supervised framework effectively captures key features relevant to downstream tasks without supervision. The alignment between embedding variance and traditional feature importance further highlights the model’s ability to learn meaningful data representations.
@@ 6 Conclusion
Overall, we have proposed a novel augmentation-free self-supervised method for representation learning that has demonstrated strong performance across both classification and regression tasks on diverse datasets. We have investigated the characteristics of the obtained representations and demonstrated that our approach is relevant as it identifies pertinent features without access to the downstream task’s target. Moreover, most methods augmented by T-JEPA outperform or match the performance of GBDT, which is often considered the go-to method for supervised tasks on tabular data. In particular, aligned with previous work that demonstrated that ResNet was a solid alternative to GBDT for tabular data (Gorishniy et al., 2021b; Zabërgja et al., 2024), we observe that ResNet+T-JEPA outperforms all competing methods on most datasets, including GBDT. Finally, we empirically uncovered a novel regularization method for transformers on tabular data by including a regularization token that prevents from entering collapsed training regimes.
@@ Acknowledgement
Appendix A Experimental settings
@@ A.3 Projection layer
Appendix B Compute and training time

### v3

@@ 4.2 Results
Table 2: Comparison of different SSL models with ResNet and MLP as downstream model. See Appendix F for more detail on experimental setting. We report and average over 20 runs and the corresponding standard deviation. The last columns displays the average rank over each dataset, the lower the better.
[…]
Downstream Models
As displayed in Table 1, T-JEPA pre-training improved the performance of all evaluated models except on the Higgs (HI) dataset where only the performance of ResNet improves when augmented by T-JEPA. Interestingly, despite their design advantages for tabular data, recent attention-based models such as FT-Transformer and AutoInt, though improved by T-JEPA, did not surpass architectures like ResNet and MLP, which exhibited the most significant gains with T-JEPA. This outcome suggests that T-JEPA may complement the inductive biases of more traditional architectures like ResNet and MLP, which, despite their simpler design, are better equipped to leverage the feature representations learned through T-JEPA. Interestingly, when compared to gradient boosted decision trees, we observe that augmented with T-JEPA representation, the deep methods regularly obtain the best performance.
Comparison to SSL methods
As displayed in Table 2, when compared to existing SSL methods, ResNet+T-JEPA obtains the lowest average rank while MLP+T-JEPA obtains the second lowest average rank. This emphasizes the relevance of T-JEPA for generating representation that foster performance on both regression and classification tasks. Other best performing alternative include MLP+SwitchTab and MLP+PTaRL that obtain respectively the third and fourth lowest average rank.
Overall these experiments demonstrate that T-JEPA contributes to increasing the performance of a wide range of downstream models. Moreover, it appears that T-JEPA outperforms existing SSL methods for tabular data when tested with MLP and ResNet as the downstream model.
@@ 5.1 Representation space evaluation
Two critical properties are considered desirable in a representation space: uniformity and alignment (Wang and Isola, 2020). Uniformity measures how well information contained in the original data space is preserved. Alignment describes a representation space in which semantically close samples are close to one another, while different samples should be distant. An edge case where both uniformity and alignment are not achieved is the one of representation collapse (see section 5.2) that describes a situation in which all samples are mapped to the same representation in the latent space.
Metrics
We propose two complementary metrics to measure whether our representations are collapsed or satisfy the uniformity and alignment properties. First, to measure distribution consistency and alignment, we rely on the Kullback-Leibler divergence ($D_{KL}(\uparrow)$): collapsed representations would imply similar distributions between diverse samples in the embedding space. Thus, lower values would indicate collapse as different samples have indistinct distributions. In contrast, higher values show that dissimilar samples tend to be located farther apart. We expect $D_{KL}$ to increase as training progresses. Second, we rely on the $\texttt{uniformity}(\uparrow)$ metric (Wang and Isola, 2020) to measure how much of the representation space is utilized. Collapsed representations would imply that samples are tightly clustered in a small region of the latent space, achieving low uniformity. On the contrary, desirable representations are more uniformly spread out, effectively using the embedding space. See Appendix G for more detail on the metrics.
Embedding space characterization
@@
To assess whether T-JEPA generates representations uniformly spanned across the embedding space, we display in Figure 2 the obtained sample representations for the JA dataset at different training epochs. We randomly sample 50,000 points and rely on the PaCMAP (Wang et al., 2021b) dimensionality reduction technique to reduce their representations in the embedding space to two dimensions for visualization purposes. We observe in Figure 2 that T-JEPA effectively utilizes the representation space, evolving from an initial collapsed distribution towards a more structured arrangement as the training progresses. This behavior is indicative of the model’s capacity to learn distinct and meaningful representations, which are crucial for downstream tasks. We also display in Table 3 how the uniformity score, KL divergence ($D_{\text{KL}}$), and the euclidean distance ($\|\cdot\|_{2}$) vary during training. At each epoch, we randomly select 50,000 sample representations from the JA dataset and compute these metrics. For both KL-Divergence and euclidean distance, we report the average pairwise KL-Divergence and euclidean distance for the 50,000 samples previously selected. The increasing KL divergence and euclidean distance demonstrate that the model effectively avoids representation collapse and better satisfies the alignment property as training progresses. Indeed, T-JEPA’s training objective pushes samples farther from one another which facilitates discrimination between samples for supervised tasks. Finally, increasing uniformity across epochs is in line with the representations displayed in Figure 2 and demonstrates that the training objective preserves information from the original dataspace.
@@ 5.2 Representation Collapse
Figure 3: Training regime of Joint-Embedding Predictive Architectures on tabular (JA) and image (ImageNet-1K) data. We display on the right a randomly selected sample’s representations for each critical part of the training process. The subfigures (a) to (d) illustrate the evolving outputs of the context encoder $h_{\text{context}}\in\mathbb{R}^{d\times h}$. In each heat-map, rows correspond to the $d$ features, while columns represent the $h$ hidden dimensions. (a) describes the initial random initialization, (b) the collapsed equilibrium, (c) the regularization effect pushing the weights outside of the collapsed equilibrium, and (d) the convergence.
@@
We observe that JEPA-based models undergo an initial representation collapse due to the EMA relation between the weights of the context and label encoders.
@@
Indeed, we retrained from scratch I-JEPA (Assran et al., 2023) on ImageNet-1K using their official implementation and hyperparameters [^1] For computational purposes, we reduced the batch size to $16$ and kept unchanged the rest of the hyperparameters. and observed a similar training regime as for T-JEPA. First, the loss collapses close to 0 after a few iterations; then, regularization starts pushing the model’s weights toward a non-trivial equilibrium. We display in Figure 3 the training regimes of both T-JEPA and I-JEPA.
@@
While other JEPA methods do not appear to require further regularization, we observe that appending a regularization token [REG] to the sample’s representations is instrumental in escaping training collapse regimes. As displayed in Figure 4, when training T-JEPA on the Jannis dataset without including any regularization token, the optimization process does not manage to escape the initial collapse. On the contrary, when including one or more tokens, the optimization process can escape this initial collapse and pushes the weights towards a non-trivial equilibrium.
@@ 5.3 Comparison with Gradient Boosted Decision Trees
To try and understand why T-JEPA enabled some approaches to match the performance of GBDT, we investigated whether high-variance features in the latent space correlate with feature importance for downstream tasks. While the $d$ representations do not exactly correspond to a one-to-one relation with the corresponding feature in the original dataspace, index embeddings allow the obtained encoded representations to still hold feature-related information.
@@
Including index embeddings does not eliminate token mixing from residual/self-attention modules, nevertheless it still helps maintain a degree of alignment between features and representations. The embedding variance was computed as the standard deviation of each feature’s values across hidden dimensions, normalized by the dimension-wise mean. We ranked features by their average embedding variance across all samples and compared these rankings to those generated by traditional supervised methods, including XGBoost and permutation importance, using Kendall’s $\tau$ for rank similarity. Overall, our analysis reveals a strong correlation between high-variance features in the T-JEPA embeddings and those identified as important by supervised methods, despite T-JEPA being trained without target labels. Figure 5 provides a detailed comparison. The embedding variance ranking ($\sigma_{\text{embed}}$) shows significant correlation with XGBoost ($\tau=0.44$, with $p\text{-value}=1.73e^{-6}$). A random generated rank is provided for comparison. These findings suggest that T-JEPA’s self-supervised framework effectively captures key features relevant to downstream tasks without supervision. The alignment between embedding variance and traditional feature importance further highlights the model’s ability to learn meaningful data representations.
@@ 6 Conclusion
Overall, we have proposed a novel augmentation-free self-supervised method for representation learning that has demonstrated strong performance across both classification and regression tasks on diverse datasets. We have investigated the characteristics of the obtained representations and demonstrated that our approach is relevant as it identifies pertinent features without access to the downstream task’s target. Moreover, most methods augmented by T-JEPA outperform or match the performance of GBDT, which is often considered the go-to method for supervised tasks on tabular data. In particular, aligned with previous work that demonstrated that ResNet was a solid alternative to GBDT for tabular data (Gorishniy et al., 2021; Zabërgja et al., 2024), we observe that ResNet+T-JEPA outperforms all competing methods on most datasets, including GBDT. Finally, we empirically uncovered a novel regularization method for transformers on tabular data by including a regularization token that prevents from entering collapsed training regimes.
@@ Acknowledgement
Appendix A T-JEPA training settings
@@ A.3 Projection layer
Appendix B Downstream models’ architecture
Similarly to T-JEPA, we employed Bayesian optimization to tune the hyperparameters of the downstream model’s architecture. We give details on each architecture and their corresponding hyperparameters in the present section.
B.1 MLP
The Multi-Layer Perceptron (MLP) architecture employed in this work is designed to effectively transform input features. The transformation begins with a linear projection of the input $\mathbf{z}$, which can either be the raw input $\mathbf{x}$ or an embedding projection, into a hidden representation of dimensionality $h$:
$\mathbf{h}^{(0)}=\mathbf{W}_{1}\mathbf{z}+\mathbf{b}_{1}.$
The model then applies the hidden layers sequentially:
$\mathbf{h}^{(l+1)}=\text{BatchNorm}\big{(}\text{Dropout}\big{(}\text{ReLU}\big{(}\mathbf{W}_{l+1}\mathbf{h}^{(l)}+\mathbf{b}_{l+1}\big{)}\big{)}\big{)},$
where $l=0,\ldots,L-1$. Here is the selected hyper-parameters for each dataset:
Table 6: Hyperparameters of MLP Model for Each Dataset
Dataset Dropout Encoder Type Learning Rate Weight Decay Hidden Layers
HE $0.5478$ linear_flatten $8.97\times 10^{-2}$ $45\times 10^{-4} 44$
HI $0.4203$ linear_per_feature $1.84\times 10^{-5}$ $7.48\times 10^{-4}$ 99
JA $0.4923$ linear_per_feature $7.31\times 10^{-4}$ $2.09\times 10^{-6}$ $13$
AD $0.2311$ linear_per_feature $1.35\times 10^{-4}$ $6.43\times 10^{-4}$ $13$
CA $0.0310$ linear_flatten $1.19\times 10^{-5}$ $1.96\times 10^{-5}$ 33
AL $0.2331$ linear_flatten $4.87\times 10^{-4}$ $1.38\times 10^{-4}$ 44
MNIST $0.3527$ linear_flatten $1.83\times 10^{-5}$ $1.47\times 10^{-4}$ 55
B.2 Improved Deep Cross Network
This section presents the enhanced DCN-V2 model architecture, designed to learn both explicit and implicit feature interactions. The code used in this work was taken from [Wang et al., 2021a]. The DCN-V2 model combines a Cross Network with a Deep Network, achieving superior expressiveness while maintaining computational efficiency. The explicit feature interactions are modeled through the Cross Network layers, defined as:
$\mathbf{x}_{l+1}=\mathbf{x}_{0}\odot(\mathbf{W}_{l}\mathbf{x}_{l}+\mathbf{b}_{l})+\mathbf{x}_{l},\quad\mathbf{W}_{l}\in\mathbb{R}^{d\times d},\;\mathbf{b}_{l}\in\mathbb{R}^{d}.$
Here, $\mathbf{x}_{0}$ represents the base features, and $\odot$ denotes element-wise multiplication. Implicit feature interactions are captured by a Deep Network, where the $l$-th layer is defined as:
$\mathbf{h}_{l+1}=f(\mathbf{W}_{l}\mathbf{h}_{l}+\mathbf{b}_{l}),\quad f(\cdot)=\text{ReLU}(\cdot).$
Table 7 summarizes the hyperparameters used in experiments.
Table 7: Hyperparameters for DCN-V2 experiments.
Dataset Cross Dropout Embedding Dim. Hidden Dim. Hidden Dropout Learning Rate Weight Decay
HE $0.0808$ $128$ $768$ $0.2926$ $8.81\times 10^{-5}$ $9.33\times 10^{-4}$
HI $0.0346$ 77 $488$ $0.0879$ $5.53\times 10^{-4}$ $3.64\times 10^{-4}$
JA $0.0702$ $70$ $386$ $0.0862$ $8.79\times 10^{-5}$ $1.21\times 10^{-7}$
AD $0.1393$ $45$ $428$ $0.2976$ $4.69\times 10^{-5}$ $1.21\times 10^{-4}$
CA $0.2578$ $66$ $704$ $0.1111$ $4.91\times 10^{-5}$ $1.11\times 10^{-5}$
AL $0.1476$ $25$ $524$ $0.0431$ $1.94\times 10^{-5}$ $1.40\times 10^{-6}$
MNIST $0.2836$ $93$ $755$ $0.0875$ $4.22\times 10^{-5}$ $1.77\times 10^{-5}$
B.3 ResNet
The ResNet model used in this work was specially tailored for tabular data. The model is formalized as follows:
$\text{ResNet}(x)$ $=\text{Prediction}(\text{ResNetBlock}(\dots(\text{ResNetBlock}(\text{Linear}(x)))))$
$\text{ResNetBlock}(x)$ $=x+\text{Dropout}(\text{Linear}(\text{Dropout}(\text{ReLU}(\text{Linear}(\text{BatchNorm}(x))))))$ (2)
$\text{Prediction}(x)$ $=\text{Linear}(\text{ReLU}(\text{BatchNorm}(x)))$
Table 8: Hyperparameters for ResNet Variants
Dataset d_block dropout1 dropout2 lr n_blocks
HE $512$ $0.459$ $0.461$ $74\times 10^{-4} 22$
AD $440$ $0.042$ $0.137$ $8.90\times 10^{-4}$ 55
CA $465$ $0.313$ $0.002$ $5.26\times 10^{-5}$ 44
HI $506$ $0.234$ $0.031$ $24\times 10^{-5} 22$
AL $476$ $0.044$ $0.049$ $1.32\times 10^{-5}$ 88
JA $355$ $0.137$ $0.005$ $4.60\times 10^{-5}$ 77
MNIST $497$ $0.265$ $0.050$ $6.25\times 10^{-4}$ 44
B.4 AutoInt
AutoInt is a neural network leveraging self-attention mechanisms for automatic feature interaction learning. The input features $\mathbf{x}\in\mathbb{R}^{n}$ are embedded into vectors $\mathbf{e}_{i}\in\mathbb{R}^{d}$ that interact through multi-head attention. These interactions produce interaction-adjusted embeddings $\tilde{\mathbf{e}}_{i}$. A residual connection ensures the retention of low-order information:
$\mathbf{e}_{i}^{\text{Res}}=\text{ReLU}(\tilde{\mathbf{e}}_{i}+\mathbf{W}_{\text{Res}}\mathbf{e}_{i}).$
The final prediction is computed as:
$\hat{y}=\sigma(\mathbf{w}^{\top}[\mathbf{e}_{1}^{\text{Res}}\oplus\dots\oplus\mathbf{e}_{M}^{\text{Res}}]+b),$
where $\sigma(x)$ represents the sigmoid function.
The hyperparameter configurations for different datasets are summarized in Table 9.
Table 9: Hyperparameter configurations for the AutoInt model across datasets.
Dataset d_token Num. layers Learning rate Residual Dropout Weight Decay
AD $200$ 11 $1.717\times 10^{-3}$ $0.060$ $1.782\times 10^{-7}$
CA $238$ 88 $2.141\times 10^{-4}$ $0.071$ $1.892\times 10^{-6}$
HI $166$ 1 $6.913\times 10^{-4}$ $0.043$ $1.385\times 10^{-7}$
AL $228$ 33 $2.280\times 10^{-3}$ $0.011$ $4.667\times 10^{-4}$
JA $32$ 66 $9.680\times 10^{-4}$ $0.090$ $1.234\times 10^{-4}$
HE $128$ 66 $5.193\times 10^{-5}$ $0.055$ $7.590\times 10^{-4}$
MNIST $252$ 22 $5.831\times 10^{-4}$ $0.091$ $5.107\times 10^{-5}$
B.5 FT-Transformer
The FT-Transformer processes tabular data using three key stages: feature tokenization, sequential Transformer layers, and a prediction head. Feature tokenization encodes each numerical feature $x_{j}$ and categorical feature $e_{j}$ as:
$T_{j}^{\text{(num)}}=b_{j}^{\text{(num)}}+x_{j}W_{j}^{\text{(num)}},\quad T_{j}^{\text{(cat)}}=b_{j}^{\text{(cat)}}+e_{j}W_{j}^{\text{(cat)}}.$
These embeddings are concatenated into a matrix $T\in\mathbb{R}^{k\times d}$, where $k$ is the number of features and $d$ is the embedding dimension.A [CLS] token is prepended to $T$, and $L$ Transformer layers are applied, iteratively updating the representation as follows:
$T_{i}=\text{MHSA}(\text{LN}(T_{i-1}))+T_{i-1},\quad T_{i}=\text{FFN}(\text{LN}(T_{i}))+T_{i},$
where LN is layer normalization. The prediction head computes the final output by processing the [CLS] token with a normalized and activated linear transformation:
$\hat{y}=W_{2}\cdot\text{ReLU}(W_{1}\cdot\text{LayerNorm}(T_{L}^{[\text{CLS}]})).$
The following table summarizes the hyperparameters for the FT-Transformer model applied to various datasets.
Table 10: Hyperparameter Configuration per Dataset
Dataset Att. Dropout Num. Heads Block Size Hidden Dim Learning Rate
HE $0.0157$ $16$ $64$ $128$ $1.08\times 10^{-3}$
HI $0.3334$ 44 $128$ $64$ $2.25\times 10^{-4}$
JA $0.0934$ $16$ $64$ $256$ $1.33\times 10^{-4}$
AD $0.3426$ $16$ $128$ $32$ $1.35\times 10^{-4}$
CA $0.0745$ 88 $128$ $512$ $1.17\times 10^{-3}$
AL $0.3246$ 88 $256$ $64$ $8.27\times 10^{-4}$
MNIST $0.4259$ 88 $32$ $256$ $7.97\times 10^{-5}$
Appendix C Compute and training time